OpenStudy (anonymous):

a rectangular prism, open at the top, has a length of 14cm, width of 9cm and the surface area is 471cm^2. find the height of the prism

OpenStudy (anonymous):

working out please..

OpenStudy (anonymous):

2lh+2bh=surf area 2h(l+b)=471 2h*23=471 h=471/(23*2)